OpenStudy (anonymous):

PLEASE HELP An object is dropped from a height of 25 meters. At what velocity will it hit the ground? A. 7.0 meters/second B. 11 meters/second C. 22 meters/second D. 49 meters/second E. 70 meters/second

OpenStudy (anonymous):

mgh = 1/2 * m v^2 25*9.8 = 1/2 v^2 v = ~22 meters/second

OpenStudy (anonymous):

Vf^2=Vi^2+2ad, where a=9.8, d=25m, and Vi=0m/s. Vf is the final velocity.